Enough of a monoprotic acid is dissolved in water to produce a 0.0178 M solution. The pH of the resulting solution is 2.36. Calculate the pK for the acid. Number
Answer #1

let HA represent monoprotic acid. HA +H2O -------->H3O+ + A-

pH= -log[H3O+] , 2.38= -log [H3O+] , [H3O+] = 10(-2.38) = 0.0042 = [A-]

[HA] at equilibrium = 0.0178-0.0042 = 0.0136

Ka= [H3O+] [A-]/ [HA] =0.0042*0.0042/ 0.0136 = 0.001297

pKa= 2.88